Suppose, you have a block of Cu of the size 1 mole. It is initially at 300 K
and is brought to a temperature of 400 K over 2 steps where temperature is
first brought to T = 350 K and then to 400K by contact with appropriate
thermal reservoirs. Treat the combination of the Cu block and the thermal
reservoir at every step to be an isolated system. Compute ∆S for the
process and compare your results with that of the previous problem. Can
you comment on the results.
1
Expert's answer
2020-12-24T11:41:25-0500

Answer

Entropy in first case

"S_1=\frac{nC_vdT}{T}=\frac{1\times0.38\times100}{300}"

"=0.126" J/k

In second case

"S_2=\frac{nC_vdT}{T}=\frac{1\times0.38\times50}{350}"

"=0.054" J/k

So change in entropy

"\Delta S=S_1-S_2=0.126-0.054\\=0.072J/k"